OpenStudy (anonymous):

A maths teacher bought 40 calculators at 8.20 dollar each and a number of other calculators costing 2.95 dollar each.in all she spent 387 dollar.how many of the cheaper calculators did she buy?

OpenStudy (anonymous):

Let the number of cheap calculators be X. Total cost of costlier calculators = 40 * 8.20 Total cost of cheaper calculators = 2.95 * X Total cost = 387 = 40*8.20 + 2.95*X Solve for X :D

OpenStudy (jhannybean):

costlier calculators lmao.

OpenStudy (anonymous):

lol whatever.. :D

OpenStudy (jhannybean):

they're called TI-89s :P

OpenStudy (anonymous):

thanks @vikrantg4

OpenStudy (jhannybean):

No reason to guess and check. in @vikrantg4's method, instead of having two equations we combined both necessities in one. The ratio of 40 calcs at 8.20 each and the "other"(x) calculators at 2.95 each that altogether totaled 387 dollars.